Gaseous butane will react with gaseous oxygen O2 to produce gaseous carbon dioxide CO2 and gaseous water H2O. Suppose 4.1 g of butane is mixed with 4.37 g of oxygen. Calculate the maximum mass of water that could be produced by the chemical reaction. Be sure your answer has the correct number of significant digits.
1
Expert's answer
2019-10-04T04:40:39-0400

      0.1366mol             Xmol

2C4H10 + 13O2 = 8CO2 + 10H2O

             13mol                10mol

n(C4H10) = m(C4H10)/M(C4H10) = 4.1g/58g/mol = 0.0707 mol (excess).

n(O2) = m(O2)/M(O2) = 4.37g/32g/mol = 0.1366 mol (deficiency).

n(H2O) = X = (0.1366 × 10 / 13) mol = 0.1051 mol.

m(H2O) = n(H2O) × M(H2O) = 0.1051 mol × 18 g/mol = 1.8918 g ≈ 1.89 g.